OpenStudy (anonymous):

Find the value by referring to the graph of the sine function. sin(-7/2 pi)

OpenStudy (anonymous):

\[\sin (-\frac{ 7 }{ 2 }\pi)\]

OpenStudy (anonymous):

is one of your answers 0.187 or 0.27

OpenStudy (anonymous):

idk? lol. it's not multiple choice

OpenStudy (anonymous):

oh than try the second because i rounded up

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

do you have a calculater???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

You can also use the unit circle :D

OpenStudy (anonymous):

its easier in my opinion Dx

OpenStudy (anonymous):

how would I do that @Nemo1223

OpenStudy (anonymous):

okay so you push the sin button and the you do 7*3.14 and then you will get 21.98 than you divide it by 2 and you get 0.1871414612

OpenStudy (anonymous):

you could just find a unit like this one http://upload.wikimedia.org/wikipedia/commons/4/4c/Unit_circle_angles_color.svg

OpenStudy (anonymous):

x is sin and y is cos

OpenStudy (anonymous):

find -7/4pi

OpenStudy (anonymous):

you would be going clock wise